Inc. has a 8.0% coupon, with semiannual payments. YTM is 8.78%. Maturity is 5 yrs. What's the price? Hint: remember for semi-annual, divide both the coupon rate and YTM by 2, and multiply the number of periods by 2. This is exactly like the semi-annual bond problems in the book and video.

$1,650.75
$968.97
$1,003.02
$1,318.23
$1,287.20
Answer #1

Information provided:

Par value= future value= $1,000

Time= 5 years*2= 10 semi-annual periods

Interest rate= 8.78%/2= 4.39% per semi-annual period

Coupon rate= 8%/2= 4%

Coupon payment= 0.04*1,000= $40

The price of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

N= 10

I/Y= 4.39

PMT= 40

Press the CPT key and PV to compute the present value.

The value obtained is 968.97.

Therefore, the price of the bond is $968.97.

Hence, the answer is option b.

In case of any query, kindly comment on the solution.